The volume of a cone is 150pi cubic inches. The height of the cone is 30 inches. What is the radius of the cone in inches?
Answer by checkley71(8403)

150PI=PI*R~2*30 OR CANCELLING OUT THE PIs WE GET 150=R~2*30 OR R~2=150/30 OR
R~2=5 OR R=SQRT5 OR R=2.2360679